OpenStudy (anonymous):

What equation is equal to (x+3)^2? A : x^2+6 B : x^2+9 C : x^2+6x+9 D : x^2+3x+9

OpenStudy (callisto):

(x+3)^2 = x^2 +2(x)(3)+(3)^2 = x^2 +6x+9

OpenStudy (anonymous):

SO whats the final answer

OpenStudy (callisto):

what do you think?

OpenStudy (anonymous):

x^2 +6x +9

OpenStudy (anonymous):

That ^ ,

OpenStudy (callisto):

yes